OpenStudy (anonymous):

A professional basketball team plays in an arena that holds 20 000 spectators. Average attendance at each game has been 14 000. The average ticket price is $75. Market research shows that for each $5 reduction in the ticket price, attendance increases by 800. Find the price that will maximize revenue.

OpenStudy (dan815):

sparkly did u see the anwer to your other qusetion?

OpenStudy (anonymous):

yes thank you

OpenStudy (anonymous):

\[ \text{population}(x) = \max(14000-75x,\; 20000) \\ \text{revenue}(x) = \text{population}(x)\cdot x =\max(14000-75x,\; 20000)x \\ \]

OpenStudy (anonymous):

i don't understand how you got those

OpenStudy (anonymous):

Hmmm I might have made a mistake.

OpenStudy (anonymous):

When \(x=75\), the population is \(14000\)

OpenStudy (anonymous):

When \(x = 70\), the population is \(14800\)

OpenStudy (anonymous):

You can use this to find the population function, since it is a line...

OpenStudy (anonymous):

This is from point-slope formula: \[ (y -14000) = -\frac{800}{5}(x-75) \]